OpenStudy (anonymous):

find the product and simplify (3s-6) / s-3 * (2s+2)/ s-2

OpenStudy (anonymous):

\[\frac{ \left( 3s-6 \right) }{ s-3 }*\frac{ \left( 2s+2 \right) }{ s-2 }\]

OpenStudy (campbell_st):

factor where possible